OpenStudy (anonymous):

Find the Cartesian equation of x=2+sint and y=2cost-1

OpenStudy (asnaseer):

get an expression for \(\sin(t)\) and \(\cos(t)\) then use the identity: \(\sin^2(t)+\cos^2(t)=1\)

OpenStudy (cwrw238):

sint = x - 2 cost = (y + 1)/2 now use the identity sin^2t + cos^t = 1

OpenStudy (anonymous):

Okay that's what I got. Thank you very much.

OpenStudy (asnaseer):

yw :)